zepdrix (zepdrix):

I need a refresher on matrix multiplication (2x2 times a 2x2). So I have these two matrices:\[\Large X^{-1}(0)\quad=\quad \left[\begin{matrix}0 & 1/3 \\ 1 & -1/3\end{matrix}\right]\]\[\Large X(2\pi)\quad=\quad \left[\begin{matrix}e^{2\pi} & e^{-2\pi} \\ 3e^{2\pi} & 0\end{matrix}\right]\] The multiplication is like this:\[\Large B\quad=\quad X^{-1}(0)\cdot X(2\pi)\] I seem to keep coming up with:\[\Large B\quad=\quad \left[\begin{matrix}0 & e^{2\pi} \\ e^{-2\pi} & 0\end{matrix}\right]\]

zepdrix (zepdrix):

But the answer is supposed to be:\[\Large B\quad=\quad \left[\begin{matrix}e^{2\pi} & 0 \\ 0 & e^{-2\pi}\end{matrix}\right]\] Heeeeeelp :( Is there some rule that allows me to switch places like that? Or am I just messing up the multiplication?

zepdrix (zepdrix):

@dan815 @satellite73

OpenStudy (dan815):